OpenStudy (anonymous):

Nakim simplified 3 sqrt 2x + x sqrt 8x – 5 sqrt 18x and got -10x sqrt 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression.

OpenStudy (anonymous):

sqrt(8x)=sqrt(2)*sqrt(2)*sqrt(2x)=2*sqrt(2x) sqrt(18x)=sqrt(3)*sqrt(3)*sqrt(2x)=3*sqrt(2x) so 3*sqrt(2x)+x*2*sqrt(2x)-5*3*sqrt(2x)=sqrt(2x)*(3+2x-15)=(2x-12)*sqrt(2x) so Nakim did wrong :)